# Control of a commercial vehicle by a tetraplegic human using a bimanual brain-computer interface

### v1

@@ Abstract
Brain-computer interfaces (BCIs) read neural signals directly from the brain to infer motor planning and execution. However, the implementation of this technology has been largely limited to laboratory settings, with few real-world applications. We developed a bimanual BCI system to drive a vehicle in both simulated and real-world environments. We demonstrate that an individual with tetraplegia, implanted with intracortical BCI electrodes in the posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC), reacts at least as fast and precisely as motor intact participants, and drives a simulated vehicle as proficiently as the same control group. This BCI participant, living in California, could also remotely drive a Ford Mustang Mach-E vehicle in Michigan. Our first teledriving task relied on cursor control for speed and steering in a closed urban test facility. However, the final BCI system added click control for full-stop braking and thus enabled bimanual cursor-and-click control for both simulated driving through a virtual town with traffic and teledriving through an obstacle course without traffic in the real world. We also demonstrate the safety and feasibility of BCI-controlled driving. This first-of-its-kind implantable BCI application not only highlights the versatility and innovative potentials of BCIs but also illuminates the promising future for the development of life-changing solutions to restore independence to those who suffer catastrophic neurological injury.
@@
This study develops an intracortical BCI system that allows a tetraplegic person to volitionally drive a commercial vehicle.
@@ INTRODUCTION
Brain computer interfaces (BCIs) are changing the ways in which humans interact with their environment. BCIs translate neural activity from the brain and can help individuals overcome motor disabilities, enabling their control of robotic limbs, wheelchairs, exoskeletons, and functional electrical stimulation of the body [1, 2, 3, 4, 5]. BCIs also facilitate communication by decoding speech articulation, interpreting motor imagery for handwriting, and enabling interactions with keyboards [6, 7]. In addition, they support the use of computers, tablets, and smartphones by controlling cursor movement, clicks, and keyboards [8, 9, 10]. For those with severe paralysis, BCI technology increases their opportunities to restore independence, return to work, and enhance social interactions [11].
In 2019, approximately 20.6 million people worldwide lived with spinal cord injury (SCI) [12], roughly 60% of whom were estimated to have tetraplegia, meaning that all four of their limbs were affected by paralysis. Tetraplegia has devastating effects on an individual’s independence, functionality, mental health, and overall quality of life. Beyond activities of daily living, individuals with SCI are often dependent on caregivers, partners, and other family members for transportation needs. Driving is frequently regarded as a key component of individual autonomy, particularly in societies such as the United States, where the ownership of personal vehicles and the necessity of driving are widespread. These capabilities are considered essential to facilitate independent mobility and access to social, economic, and healthcare resources. The capacity of personally controlling a vehicle not only encompasses transportation but also preserves a sense of agency and independence, and its loss constitutes a decline in functional independence and autonomy.
@@
In this study, we present the first intracortical BCI system that allows real-time bimanual control of a commercial vehicle by a tetraplegic human participant (see Figure 1). A multidisciplinary group spanning academia and industry built a BCI driving system that allowed a tetraplegic BCI participant, BCI-JJ, to remotely operate a Mustang Mach-E (Ford Motor Company, Dearborn, MI) in closed test environments without other traffic or pedestrians to ensure appropriate safety control. As part of an FDA-approved clinical trial (ClinicalTrials.gov number NCT01958086), we implanted BCI-JJ intracortical NeuroPort Electrodes (Blackrock Neurotech, Salt Lake City, UT) in both the posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC) (see figure S1). BCI-JJ employed mental imagery of both upper limbs for control, which led to a more intuitive separation of driving functions across the left and right limbs. Using this bimanual control scheme, BCI-JJ could volitionally control the vehicle for both extended navigation in the real world and more specific tests of driving capabilities. In experimentally constrained laboratory tests, BCI-JJ, using our BCI system, reacted at least as fast and precisely as a motor intact control group performing the same tasks with a standard computer mouse.
Figure 1: The BCI control system diagram for teledriving a Ford Mustang Mach-E vehicle. Our setup included the decoder and display computers at the BCI test site in California and the in-vehicle controller computer at the vehicle test site in Michigan. Brain signals were recorded from BCI-$J$’s posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC) via the intracortical NeuroPort Electrodes (Blackrock Neurotech, Salt Lake City, UT) and processed by the NeuroPort system (Blackrock Neurotech, Salt Lake City, UT). The decoder computer extracted the neural features from the acquired signals, decoded the motor intention, and generated a corresponding motor command. This motor command was transmitted to the display computer and then sent to the in-vehicle controller via TCP and UDP. At the vehicle test site, the in-vehicle controller computer executed the motor command for the vehicle. This computer also recorded a live-stream video from a camera mounted on the vehicle. The video feedback and the vehicle state were transferred to the display computer at the BCI test site for BCI-JJ to watch and respond in real-time.
This work represents both a BCI-enabled driving system with the most advanced capability to date and a model for meaningful academic-industry collaboration in the rapidly growing field of BCIs. Building this system required neuroscientific development to provide novel bimanual capabilities enabled by implants across the cortical motor system, as well as significant technology development to allow remote moment-by-moment control of a full-sized commercial vehicle.
@@ RESULTS
A
B
Figure 2: Detailed flowcharts of the BCI-enabled driving system. (A) The detailed mechanism of BCI neural signal decoding and driving control of either a Mach-E commercial vehicle remotely or a CARLA simulated vehicle. (B) The vehicle control architecture of a Ford Mach-E.
@@
Simple
Reaction
Time
Unity 2018.4.23f1
Simple reaction time &\&
performance measures
Yes
Braking
Reaction
Time
CARLA 0.9.13 ++
brake overlay
Braking reaction time &\&
performance measures
Yes
Mcity
Teledriving
Ford Mach-E ++
speed and steering
overlay
4 random routes (with
turnings, stops, lane
switches, roundabouts)
No
Simulated
Town
Driving
CARLA Lb 2.0 ++
speed, steering
and brake overlay
A fixed route (with
traffic lights, turnings,
lane switches)
Yes
Obstacle
-course
Teledriving
Ford Mach-E ++
speed, steering
and brake overlay
4 fixed routes (with
stops, turnings, lane
switches, roundabouts)
No
To evaluate the capability of our BCI system to control a vehicle, we conducted a series of five tasks, progressing from in-lab to more complex, real-world applications (see Table 1). First, we compared the speed and proficiency of BCI-JJ to a control group of motor intact individuals during two reaction time tasks. Task 1 tested the reaction times in a simple task [35, 36], in which each participant made an attempted or actual click, via our BCI click decoder (see the “BCI decoder system” section) or a computer mouse, in response to a target appearing on a computer screen (see Figure 3A). Task 2 tested the braking reaction times during simulated driving in CARLA 0.9.13 [37] in which each participant was asked only to brake (again via a click) in response to each random obstacle appearing in front of a virtual vehicle driving on a predetermined course with random settings of sunlight and shadows (see Figure 3B). We evaluated the reaction times of the motor intact control group compared to BCI-JJ in both reaction time tasks, finding that BCI-JJ had similar reaction performance measures and significantly faster reaction times than most of the motor intact participants.
@@
Figure 3: Display of five tasks and the driving routes. (A) Simple reaction time task. (B) Simulated braking reaction time task in CARLA 0.9.13. (C) BCI-controlled teledriving of a Mach-E without traffic in the Mcity test facility. (D) Four teledriving routes in Mcity. (E) Simulated town driving with traffic in the CARLA Leaderboard 2.0. (F) One simulated driving route in Town 12 of CARLA. (G) BCI-controlled teledriving of a Mach-E without traffic on an obstacle course. (H) Four teledriving route options on the loop-shaped obstacle course.
As a proof-of-concept, we then tested the ability of our BCI system to remotely control the speed and steering of a Ford Mustang Mach-E vehicle in real time (see Figures 1 and 2A). Using a cursor-movement-based BCI decoder (see the “BCI decoder system” section) and control system, BCI-JJ navigated the Mach-E vehicle in Mcity, a closed test facility in Michigan made to represent an urban driving environment without traffic, in Task 3 (see Figures 3C and 3D). As shown in Figure 2A, the overlay module obtained the decoded x and y cursor movement values from the effector controller module and displayed them as positions of the blue and red circles, which would be post-processed for steering and speed control of the vehicle, respectively (see the “BCI-enabled driving” section for a full description).
@@
Once the basic safety and feasibility of vehicle control was established, we implemented a more complex simulated driving task modified from the CARLA Leaderboard 2.0, using the bimanual cursor-and-click control, to evaluate our BCI control system in more realistic traffic situations. In Task 4, which mimicked the downtown driving scenarios in a big city with skyscrapers and busy traffic (see Figures 3E and 3F), we compared the simulated driving performance of BCI-JJ to that of the motor intact control group, both with bimanual control which included click movement with the left index finger for full-stop braking and cursor movement with the right thumb for speed and steering adjustments. Like the display for the unilateral cursor control previously in Task 3, we still used the same overlay module to display bimanual control, with each click-to-brake information represented by an immediate decrease-to-the-bottom of the vertical position for the red circle, which would be post-processed as an immediate speed decrease to zero of the vehicle (see Figure 2A and the “BCI-enabled driving” section for a full description). Using simulated driving proficiency scores, calculated based on the occurrence of traffic infractions and the completeness of the route in Task 4, we were able to show that BCI-JJ had similar driving proficiency compared to the motor intact control group, indicating that the BCI control system can achieve essential functions of a vehicle while driving.
Finally, to test the translatability of BCI vehicle control, we used the bimanual BCI control system described in the previous simulated driving task (i.e., Task 4) to control the same Ford Mustang Mach-E vehicle as in the Mcity teledriving task (i.e., Task 3), to remotely navigate an obstacle course in real time in Task 5 (see Figures 3G and 3H). This course included components of a typical driving test for a motor intact person (e.g., full stops, lane switches, turnings, roundabouts, and obstacle avoidance). BCI-JJ was able to fully navigate the Mach-E vehicle through the obstacle course with a high degree of proficiency, demonstrating the first implementation of bimanual intracortical BCI control of a full-sized commercial vehicle in a real-world driving environment.
@@
The main purpose of our study is to robustly apply our BCI system to real-world applications beyond the laboratory. The prerequisite for achieving this aim is to prove, in a controlled laboratory environment, that a participant using our BCI system can react as fast and precisely as individuals with intact motor functions, while controlling potential biases related to age and gender. Therefore, we first designed two reaction time tasks with various target and scenario complexities but the same concept of responding with a decoded or actual click as soon as a target appeared. The simple reaction time task had a target presented at varying time intervals on an otherwise black computer screen, and the participant was asked to click with either our BCI click decoder or a computer mouse as quickly as possible each time the target appeared in each GO trial (see Figure 3A). The braking reaction time task had obstacles appeared at random on a predetermined path in the CARLA 0.9.13 driving simulator [37], and the participant was asked to click for full-stop braking of an otherwise autonomous driving vehicle whenever an obstacle appeared in its path in the GO phase (see Figure 3B). The “Reaction time tests” section has a full description of the experimental settings for the two reaction time tasks.
@@
BCI simple reaction performance among different hand effectors
To assess the basic reaction times of our BCI control system, we conducted a simple reaction time task in which BCI-JJ attempted to click with one of six different effectors: left and right index fingers, left and right ring fingers, and left and right power grips to complete this action. Decoded click movement intentions for each effector were digitally executed as mouse clicks to complete the task. BCI-JJ was asked to complete 10 runs consisting of 40 GO trials and 10 randomly interleaved NO-GO catch trials (indicated by an auditory stimulus during the ITI phase) per run. BCI-JJ repeated these 10 runs for each effector separately.
[…]
@@
Figure 4: BCI trial-based simple and braking reaction time task results among different hand effectors of BCI-$J$. (A and B) For the simple reaction time task, we collected 10 runs for each of BCI-JJ’s six effectors (i.e., right and left index fingers, ring fingers, and power grips) consisting of 40 GO trials and 10 randomly interleaved NO-GO catch trials per run. We compared simple reaction times within the valid range from 50 ms to 1000 ms in GO trials/phases among the six effectors. We also conducted comparisons of simple reaction performance measures (i.e., accuracy, sensitivity, specificity) among these effectors. (C and D) For the braking reaction time task, we collected 10 runs per right or left index finger consisting of 40 trials per run, with one NO-GO phase and one GO phase in each trial. We compared braking valid reaction times and reaction performance measures between these two effectors. Box charts whose shaded notches around the median lines do not overlap have different medians at the corrected $5\%$ significance level.
@@
We compared the valid braking reaction times, accuracy, sensitivity, and specificity between the left and right index finger effectors via two-sample t-tests for BCI-JJ (see tables S3 and S4). Like the simple reaction time task, BCI-JJ had significantly faster reaction times using the right index finger, compared to the left (i.e., 290 ms (64.9 ms SD) versus 338 ms (82.2 ms SD), with $p<0.05$) (see Figure 4C). BCI-JJ showed greater than $90\%$ accuracy, sensitivity, and specificity across both effectors throughout all trials for this simulated braking task (see Figure 4D). Comparing each performance measure (i.e., accuracy, sensitivity, and specificity) of BCI-JJ’s right and left index fingers, their differences are either at the edge of the corrected $5\%$ significance level or clearly insignificant (see table S4).
@@
Each of the 20 motor intact control participants completed 4 runs of the simple reaction time task and 5 runs of the braking reaction time task, both with the same number of trials per run and layout as the same tasks for BCI-P2.
@@
A
B
Figure 6: Trial-based valid reaction times for the simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We conducted pairwise comparisons using a Bonferroni post hoc test following a one-way ANOVA with 21 participants. For the participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of the valid reaction times between 50 ms and 1000 ms in GO trials/phases among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Among BCI-JJ and all 20 motor intact participants (M01–M20) individually, we also compared the reaction performance using a one-way ANOVA test followed by pairwise comparisons (see Figure 6 for the valid reaction times and figures S2S3S4 for the reaction accuracy, sensitivity, and specificity). During the simple reaction time task, BCI-JJ had significantly faster valid reaction times than 12 out of 20 motor intact participants, and significantly lower accuracy and sensitivity compared to 7 motor intact participants ($p<0.05$, Bonferroni corrected), which were not necessarily the same participants between metrics. There were no significant differences in sensitivity between BCI-JJ and any motor intact participants (see tables S5 and S7). During the braking reaction time task, BCI-JJ had significantly faster valid reaction times than 18 out of 20 motor intact individuals, and significantly higher sensitivity compared to 1 motor intact participant ($p<0.05$ Bonferroni corrected). There were no significant differences in accuracy and specificity between BCI-JJ and any motor intact participants (see tables S8 and S10). This indicates that BCI-JJ, via our BCI click control, reacted at least as fast and precisely as most of the motor intact control participants, no matter of the complexities of the two reaction time tasks.

### v2

# Control of a commercially available vehicle by a tetraplegic human using a brain-computer interface

@@ Abstract
Brain-computer interfaces (BCIs) read neural signals directly from the brain to infer motor planning and execution. However, the implementation of this technology has been largely limited to laboratory settings, with few real-world applications. We developed a BCI system to drive a vehicle in both simulated and real-world environments. We demonstrate that an individual with tetraplegia, implanted with intracortical BCI electrodes in the posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC), reacts at least as fast and precisely as motor intact participants. This BCI participant, living in California, could also remotely drive a Ford Mustang Mach-E vehicle in Michigan. Our teledriving tasks relied on cursor movement control for speed and steering in a closed urban test facility and through a predefined obstacle course. These two tasks serve as a proof-of-concept that takes into account the safety and feasibility of BCI-controlled driving. The final BCI system added click control for full-stop braking and thus enabled bimanual cursor-and-click control for simulated town driving with the same proficiency level as the motor intact control group through a virtual town with traffic. This first-of-its-kind implantable BCI application not only highlights the versatility and innovative potentials of BCIs but also illuminates the promising future for the development of life-changing solutions to improve independent mobility for those who suffer catastrophic neurological injury.
@@
This study develops an intracortical BCI system that allows a tetraplegic person to volitionally drive a vehicle.
@@ INTRODUCTION
Brain computer interfaces (BCIs) are changing the ways in which humans interact with their environment. BCIs translate neural activity from the brain and can help individuals overcome motor disabilities, enabling their control of robotic limbs, wheelchairs, exoskeletons, and functional electrical stimulation of the body [1, 2, 3, 4, 5]. BCIs also facilitate communication by decoding speech articulation, interpreting motor imagery for handwriting, and enabling interactions with keyboards [6, 7]. In addition, they support the use of computers, tablets, and smartphones by controlling cursor movement, clicks, and keyboards [8, 9, 10]. For those with severe paralysis, BCI technology increases their opportunities to improve independent mobility, return to work, and enhance social interactions [11].
In 2019, approximately 20.6 million people worldwide lived with spinal cord injury (SCI) [12], roughly 60% of whom were estimated to have tetraplegia, meaning that all four of their limbs were affected by paralysis. Tetraplegia has devastating effects on an individual’s independence, functionality, mental health, and overall quality of life. Beyond activities of daily living, individuals with SCI are often dependent on caregivers, partners, and other family members for transportation needs. Driving is frequently regarded as a key component of individual autonomy, particularly in societies such as the United States, where the ownership of personal vehicles and the necessity of driving are widespread. These capabilities facilitate independent mobility and access to social, economic, and healthcare resources. The capacity of personally controlling a vehicle not only encompasses transportation but also preserves a sense of agency and independence, and its loss constitutes a decline in functional independence and autonomy.
Figure 1: The BCI control system diagram for teledriving a Ford Mustang Mach-E vehicle. Our setup included the decoder and display computers at the BCI test site in California and the in-vehicle controller computer at the vehicle test site in Michigan. Brain signals were recorded from BCI-$J$’s posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC) via the intracortical NeuroPort Electrodes © Blackrock Neurotech and processed by the NeuroPort system © Blackrock Neurotech. The decoder computer extracted the neural features from the acquired signals, decoded the motor intention, and generated a corresponding motor command. This motor command was transmitted to the display computer and then sent to the in-vehicle controller via TCP and UDP. At the vehicle test site, the in-vehicle controller computer executed the motor command for the vehicle. This computer also recorded a live-stream video from a camera mounted on the vehicle. The video feedback and the vehicle state were transferred to the display computer at the BCI test site for BCI-JJ to watch and respond in real-time.
@@
In this study, we present the first-of-its-kind intracortical BCI system that allows remote driving control of a commercially available vehicle by a tetraplegic human participant (see Figure 1). A multidisciplinary group spanning academia and industry built a BCI driving system that allowed a tetraplegic BCI participant, BCI-JJ, to remotely operate a Mustang Mach-E (Ford Motor Company, Dearborn, MI) in closed test environments. Two real-world driving tasks were conducted in a closed urban test facility and on an obstacle course without other traffic or pedestrians, with controlled speed limits, predefined safety boundaries, and emergency procedures in place, to ensure appropriate safety control. As part of an FDA-approved clinical trial (ClinicalTrials.gov number NCT01958086), we implanted BCI-JJ intracortical NeuroPort Electrodes (Blackrock Neurotech, Salt Lake City, UT) in both the posterior parietal cortex (PPC) and the hand knob region of the motor cortex (MC) (see figure S1). For the teledriving tasks with the Mach-E vehicle, BCI-JJ employed mental imagery of the right thumb for cursor-movement control of speed and steering. By employing additional mental imagery of the left index finger for click control of full-stop braking, we further upgraded our BCI driving system with a more intuitive separation of driving functions across the left and right limbs. Using this bimanual control scheme, BCI-JJ could volitionally drive a virtual vehicle with the same proficiency level as a motor intact control group through a designated route in a simulated town driving environment with heavy traffic and other complex traffic elements. In experimentally constrained laboratory tests, our BCI system enabled BCI-JJ to react at least as fast and precisely as the motor intact control group performing the same reaction-time tasks with a standard computer mouse.
A
B
Figure 2: Detailed flowcharts of the BCI-enabled driving system. (A) The detailed mechanism of BCI neural signal decoding and driving control of either a Ford Mustang Mach-E vehicle remotely or a CARLA simulated vehicle. (B) The vehicle control architecture of a Ford Mach-E vehicle.
This work represents both a BCI-enabled driving system with the most advanced capability to date and a model for meaningful academic-industry collaboration in the rapidly growing field of BCIs. Building this system required neuroscientific development to provide reliable cursor movement decoding and click classification enabled by implants across the cortical motor system, as well as significant technology development to allow moment-by-moment BCI control of a full-sized commercially available vehicle remotely as a proof-of-concept and of a virtual vehicle in a simulated town with the same proficiency level as motor intact behavior.
@@ RESULTS
Simple Reaction Time Unity 2018.4.23f1 Simple reaction time &\& performance measures [Uncaptioned image] Yes
Braking Reaction Time CARLA 0.9.13 ++ braking overlay Braking reaction time &\& performance measures [Uncaptioned image] Yes
Mcity Teledriving Ford Mach-E ++ speed and steering overlay 4 random routes (with turnings, stops, lane switches, roundabouts) [Uncaptioned image] No
Obstacle -course Teledriving Ford Mach-E ++ speed and steering overlay 4 fixed routes (with stops, turnings, lane switches, roundabouts) [Uncaptioned image] No
Simulated Town Driving CARLA Lb 2.0 ++ speed, steering and braking overlay A fixed route (with traffic lights, turnings, lane switches) [Uncaptioned image] Yes
@@
Figure 3: Display of five tasks and the driving routes. (A) Simple reaction time task. (B) Simulated braking reaction time task in CARLA 0.9.13. (C) BCI-controlled teledriving of a Mach-E without traffic in the Mcity test facility. (D) Four teledriving routes in Mcity. (E) BCI-controlled teledriving of a Mach-E without traffic on an obstacle course. (F) Four teledriving route options on the loop-shaped obstacle course. (G) Simulated town driving with traffic in the CARLA Leaderboard 2.0. (H) One simulated driving route in Town 12 of CARLA.
To evaluate the capability of our BCI system to control a vehicle, we conducted a series of five tasks, progressing from in-lab to more complex and practical applications (see Table 1). First, we compared the speed and proficiency of BCI-JJ to a control group of motor intact individuals during two reaction time tasks. Task 1 tested the reaction times in a simple task [35, 36], in which each participant made an attempted or actual click, via our BCI click decoder (see the “Click classification” section) or a computer mouse, in response to a target appearing on a computer screen (see Figure 3A). Task 2 tested the braking reaction times during simulated semi-autonomous driving in CARLA 0.9.13 [37] in which each participant was asked only to brake (again via a click) the virtual vehicle in response to each random obstacle appearing in front of it driving in a virtual town environment with random settings of sunlight and shadows (see Figure 3B). We evaluated the reaction times of the motor intact control group compared to BCI-JJ in both reaction time tasks, finding that BCI-JJ had similar reaction performance measures and significantly faster reaction times than most of the motor intact participants.
As a proof-of-concept, we then tested the ability of our BCI system to remotely control the speed and steering of a Ford Mustang Mach-E vehicle in real time (see Figures 1 and 2A). Using a single-effector cursor-movement-based BCI decoder (see the “Cursor movement decoding” section) and control system, BCI-JJ navigated the Mach-E vehicle in Mcity, as free driving through randomly mapped routes without following any traffic lights or stop signs, in Task 3 (see Figures 3C and 3D) and on a predefined obstacle course, as testing many components of a standard driving task, in Task 4 (see Figures 3E and 3F). We set the maximum speed of the Ford Mach-E vehicle to 4 mph. Such low-speed vehicle control could ensure safety of the commercially available vehicle and safety drivers, stable control on a real-world obstacle course with narrow lanes, and smooth display of the video feedback. We demonstrated the first implementation of intracortical BCI control of a full-sized commercially available vehicle in a real-world driving environment.
@@
Once the basic safety and feasibility of vehicle control was established, we implemented a more complex simulated town driving task modified from the CARLA Leaderboard 2.0, using the bimanual cursor-and-click control, to evaluate our BCI control system in more realistic traffic situations. Task 5 mimicked the downtown driving scenarios in a big city with skyscrapers and heavy traffic (see Figures 3G and 3H). Similar to the low-speed teledriving tasks, we kept the maximum speed of the virtual agent vehicle at 5 mph, the same as the traffic flow speed, for more stable control and smoother display of the video feed. We compared the simulated driving performance of BCI-JJ to that of the motor intact control group, both with bimanual control which included click movement with the left index finger for full-stop braking and cursor movement with the right thumb for speed and steering adjustments. Using simulated driving proficiency scores, calculated based on the occurrence of traffic infractions and the completeness of the route in Task 5, we were able to show that BCI-JJ had similar driving proficiency compared to the motor intact control group, indicating that the BCI control system can achieve essential functions of a vehicle while driving.
@@
The main purpose of our study is to apply our BCI system to real-world applications beyond the laboratory as a proof-of-concept. The prerequisite for achieving this aim is to prove, in a controlled laboratory environment, that a participant using our BCI system can react as fast and precisely as individuals with intact motor functions, while controlling potential biases related to age and gender. Therefore, we first designed two reaction time tasks with various target and scenario complexities, using the same concept of responding with a decoded or actual click as soon as a target appeared. The simple reaction time task had a target presented at varying time intervals on an otherwise black computer screen, and the participant was asked to click with either our BCI click decoder or a computer mouse as quickly as possible each time the target appeared in each GO trial (see Figure 3A). The braking reaction time task had obstacles appeared at random on a predetermined path in the CARLA 0.9.13 driving simulator [37], and the participant was asked to click for full-stop braking of an otherwise autonomous driving vehicle whenever an obstacle appeared in its path in the GO phase (see Figure 3B). The “Reaction time tests” section has a full description of the experimental settings for the two reaction time tasks.
@@
Figure 4: BCI trial-based simple and braking reaction time task results among different hand effectors of BCI-$J$. (A and B) For the simple reaction time task, we collected 10 runs for each of BCI-JJ’s six effectors (i.e., right and left index fingers, ring fingers, and power grips) consisting of 40 GO trials and 10 randomly interleaved NO-GO catch trials per run. We compared simple reaction times within the valid range from 50 ms to 1000 ms in GO trials/phases among the six effectors. We also conducted comparisons of simple reaction performance measures (i.e., accuracy, sensitivity, specificity) among these effectors. (C and D) For the braking reaction time task, we collected 10 runs per right or left index finger consisting of 40 trials per run, with one NO-GO phase and one GO phase in each trial. We compared valid braking reaction times and reaction performance measures between these two effectors. Box charts whose shaded notches around the median lines do not overlap have different medians at the corrected $5\%$ significance level.
BCI simple reaction performance among different hand effectors
To assess the basic reaction times of our BCI control system, we conducted a simple reaction time task in which BCI-JJ attempted to click with one of six different effectors: left and right index fingers, left and right ring fingers, and left and right power grips to complete this action. Decoded click movement intentions for each effector were digitally executed as mouse clicks to complete the task. BCI-JJ was asked to complete 10 runs consisting of 40 GO trials and 10 randomly interleaved NO-GO catch trials (indicated by a low-pitch auditory stimulus during the ITI phase) per run. BCI-JJ repeated these 10 runs for each effector separately.
[…]
@@
We compared the valid braking reaction times, accuracy, sensitivity, and specificity between the left and right index finger effectors via two-sample t-tests for BCI-JJ (see tables S3 and S4). Like the simple reaction time task, BCI-JJ had significantly faster reaction times using the right index finger, compared to the left, showing 290 ms (64.9 ms SD) versus 338 ms (82.2 ms SD) with $p<0.05$ (see Figure 4C). BCI-JJ showed greater than $90\%$ accuracy, sensitivity, and specificity across both effectors throughout all trials for this simulated braking task (see Figure 4D). Comparing each performance measure (i.e., accuracy, sensitivity, and specificity) of BCI-JJ’s right and left index fingers, their differences are either at the edge of the corrected $5\%$ significance level or clearly not significant (see table S4).
@@
Each of the 20 motor intact control participants completed 4 runs of the simple reaction time task and 5 runs of the braking reaction time task, both with the same number of trials per run and layout as the same tasks for BCI-JJ.
@@
Among BCI-JJ and all 20 motor intact participants (M01–M20) individually, we also compared the reaction performance using a one-way ANOVA test followed by pairwise comparisons (see figure S2 for the valid reaction times and figures S3S4S5 for the reaction accuracy, sensitivity, and specificity). During the simple reaction time task, BCI-JJ had significantly faster valid reaction times than 12 out of 20 motor intact participants, and significantly lower accuracy and specificity compared to 7 motor intact participants ($p<0.05$, Bonferroni corrected), which were not necessarily the same participants between metrics. There were no significant differences in sensitivity between BCI-JJ and any motor intact participants (see tables S5 and S7). During the braking reaction time task, BCI-JJ had significantly faster valid reaction times than 18 out of 20 motor intact individuals, and significantly higher sensitivity compared to 1 motor intact participant ($p<0.05$, Bonferroni corrected). There were no significant differences in accuracy and specificity between BCI-JJ and any motor intact participants (see tables S8 and S10). This indicates that BCI-JJ, via our BCI click control, reacted at least as fast and precisely as most of the motor intact control participants, no matter of the complexities of the two reaction time tasks.
BCI driving performance
Via the two reaction time tasks, we established that the reaction performance of BCI-JJ with index finger clicks decoded by our BCI system was at least as fast and reliable as the average performance of the motor intact participants clicking a computer mouse. In order to verify the robustness of our BCI system in dealing with more realistic applications, such as driving a vehicle, we outfitted a Ford Mustang Mach-E vehicle to be remotely controllable in real-world environments without traffic. Via teledriving, we first tested the speed and steering control with cursor movement decoded by our BCI system in a free-driving scenario without following any traffic lights or stop signs inside an urban test facility called Mcity (see Figure 3C). Mcity is a closed test facility in Michigan made to represent an urban driving environment without traffic, whereas the obstacle course included components of a typical driving test for a motor intact person (e.g., full stops, lane switches, turnings, roundabouts, and obstacle avoidance). As shown in Figure 2A, the overlay module obtained the decoded x and y cursor movement values from the effector controller module and displayed them as positions of the blue and red circles, which would be post-processed for steering and speed control of the vehicle, respectively (see the “BCI-enabled driving” section for a full description). Then we designed an obstacle course and applied teledriving with the same Mach-E vehicle which could verify the same BCI driving system in a more applicable real-world setting mimicking many components of a standard driving test (see Figure 3E). Finally, we increased the complexity of the BCI decoder from the single-effector cursor control for steering and speed to the bimanual cursor-and-click control, which enabled more precise full-stop braking via clicks with the left index finger effector and continuous steering and speed adjustments via the horizontal and vertical cursor movement, respectively, with the right thumb. To ensure the safety, precision, and complexity of our driving evaluations with bimanual control, we temporarily switched back to a laboratory setting and compared the driving proficiency of BCI-JJ to the motor intact control group using a virtual vehicle in a simulated town environment with heavy traffic and a more sophisticated route modified from the CARLA Leaderboard 2.0 (see Figure 3G). The “BCI-enabled driving” section has full descriptions of the experimental settings for the three driving tasks.
To ensure safe driving in accordance to many components of a standard driving test, we set up safety criteria (see the “Safety criteria for driving tasks” section) and evaluated the following infractions, including the proportion of the completed route distance ($C$), the number of collisions ($N_{c}$), the number of lane deviations ($N_{l}$), and the number of running red traffic lights or stop signs if available ($N_{s}$). The infraction metric aggregated all these infractions using the following equation:
For both teledriving tasks in Mcity and through the obstacle course, BCI teledriving infractions were evaluated based on recorded videos by three independent human evaluators from the California Institute of Technology, Blackrock Neurotech, and Ford Motor Company. Please note that if the vehicle was about to collide with some dangerous obstacle, $N_{c}$ would increase by 1, and the safety driver in the test vehicle would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. If the vehicle was about to turn left at an intersection where a left-turn lane was available, it should merge to that left-turn lane before its left turn during BCI teledriving; otherwise, $N_{l}$ would increase by 1. If an evaluator was uncertain about the occurrence of any infraction component ($N_{c}$, $N_{l}$ or $N_{s}$) (e.g., whether there was a lane deviation at $t=45$ seconds in Run 2), he/she would mark it with a quantity increase of 0.5 instead of 1 (e.g., $N_{l}\mathrel{+}=0.5$).